# The near horizon dynamics of three-dimensional Einstein gravity

### v1

@@ 1 Introduction
The black hole near horizon physics has recently gained new attention in the framework of Jackiw-Teitelboim (JT) gravity Jackiw:1984je ; Teitelboim:1983ux . JT gravity as a two-dimensional dilaton gravity describes the dynamics close to the horizon of near-extremal black holes. Under appropriate boundary conditions it provides a toy model for holography in the Rindler AdS2 background, sometimes referred to as near AdS2/near CFT1 correspondence Kitaev:15ur ; Kitaev:2017awl ; Almheiri:2014cka ; Maldacena:2016hyu ; Maldacena:2016upp ; Engelsoy:2016xyb ; Cvetic:2016eiv . Its holographic description is in terms of a one-dimensional Schwarzian action at the boundary for the Euclidean time reparametrization (which is the element of the asymptotic symmetry group). Similar holographic correspondences have been scrutinized in the 2D flat Rindler background (a near horizon description of non-extremal black holes) for flat JT-models Callan:1992rs ; Cangemi:1992bj ; Afshar:2019axx ; Godet:2020xpk ; Afshar:2020dth ; Godet:2021cdl ; Afshar:2021qvi ; Afshar:2022mkf . The Euclidean boundary dynamics in all such cases is in terms of an effective 1D Schwarzian-like action for elements of the asymptotic symmetry groups Afshar:2019tvp ; Afshar:2021qvi .
@@
The structure of the paper is as follows, in section 2 after introducing the 3D Rindler gravity theory we offer a WZW formulation for the theory using angular quantization of the Chern-Simons formulation. Then via Gauss decomposition of the gauge group we obtain a liouville-like boundary action for the flat Rindler boundary condition. In section 3 we promote this theory to a reparametrization theory in terms of elements of the warped Virasoro-group. In section 4 we develop the same discipline for Rindler AdS3 and finally we discuss various aspects of the holonomy conditions in section 5.
@@ 2 Rindler boundary conditions
We are fully off-shell unless stated explicitly in the rest of the paper. The asymptotic form of the metric (2.1) is preserved by diffeomorphisms generated by a vector field $\xi$ producing a Lie derivative such that $\mathcal{L}_{\xi}g_{xr}={\mathcal{L}}_{\xi}g_{xx}={\mathcal{L}}_{\xi}g_{rr}={\mathcal{L}}_{\xi}g_{ru}=0$;
$\xi^{u}=\epsilon(u)+\mathcal{O}(1/r)\,,\qquad\xi^{x}=\alpha(u)+\mathcal{O}(1/r)\,,\quad\xi^{r}=-r\dot{\epsilon}(u)+\mathcal{O}(1)\,.$ (2.4)
@@ 2.1 Chern-Simons formulation
The action that gives 3D Einstein gravity is the Chern-Simons action up to a possible boundary term, for the gauge group ISO$(2,1)$, SO(2,2)22(2,2)( 2 , 2 ) or SO$(3,1)$ depending on the value of the cosmological constant $\Lambda=\mp\frac{1}{\ell^{2}}$ Achucarro:1986vz ; Witten:1988hc ,
@@
where $L_{\pm 1}=J_{0}\pm J_{1}$, $M_{\pm 1}=P_{0}\pm P_{1}$ and $L_{0}=J_{2}$ and $M_{0}=P_{2}$ and thus $\langle L_{1}\,M_{-1}\rangle=\langle L_{-1}\,M_{1}\rangle=-2\langle L_{0}\,M_{0}\rangle=-1$. We can recast our boundary conditions (2.1) in the first order form in terms of an algebra-valued connection, [^1] We have made a finite gauge transformation with the SO(2,2)22(2,2)( 2 , 2 ) group element $g=e^{\frac{r}{2}(M_{-1}-\frac{1}{\ell}L_{-1})}$ to eliminate the $r$-dependence. All results will be valid in the flat limit $\ell\to\infty$ for the theory with gauge group ISO$(2,1)$.
@@
$\delta I_{\text{{\tiny CS}}}=-\frac{k}{4\pi}\int\operatorname{d}\!^{3}x\,\epsilon^{\mu\nu\rho}\partial_{\mu}\langle A_{\nu}\delta A_{\rho}\rangle+\delta I_{\text{{\tiny bdy}}}\,.$ (2.9)
@@
$\langle A_{u}A_{x}\rangle=\frac{a}{2}+\mathcal{O}(1/\ell)\,.$ (2.11)
@@ 2.2 From Chern-Simons to Wess-Zumino-Witten
where $x^{\alpha}=(r,u)$ are coordinates on the $x$-constant slices, the Chern-Simons action becomes,
$I_{\text{\tiny{CS}}}=\frac{k}{4\pi}\int\operatorname{d}\!x\int\operatorname{d}\!^{2}x\,\epsilon^{\alpha\beta}\left\langle-A_{\alpha}\partial_{x}A_{\beta}+A_{x}F_{\alpha\beta}\right\rangle+I_{\text{\tiny bdy}}\,.$ (2.13)
Here by assuming that there are no nontrivial holonomies on the $x$-constant plane, a pure gauge configuration solves the constraint $F_{ru}=0$ by,
$A_{\alpha}=G^{-1}\partial_{\alpha}G\,,\qquad\alpha=r,u\,.$ (2.14)
@@
$I_{\text{\tiny CS}}$ $=\frac{k}{4\pi}\int\operatorname{d}\!^{3}x\,\langle A_{u}A_{r}^{\prime}-A_{r}A_{u}^{\prime}\rangle+I_{\text{\tiny bdy}}$
$=\frac{k}{4\pi}\int\operatorname{d}\!^{3}x\,\langle-b^{-1}Ab^{\prime}b^{-1}\partial_{r}b+b^{-1}A\partial_{r}b^{\prime}+\partial_{r}bb^{-1}b^{\prime}b^{-1}A-\partial_{r}bb^{-1}A^{\prime}-b^{-1}\partial_{r}bb^{-1}Ab^{\prime}\rangle+I_{\text{\tiny bdy}}\,.$ (2.17)
@@
$I_{\text{\tiny CS}}$ $=\frac{k}{4\pi}\left(-\Gamma[G]+\int\operatorname{d}\!^{3}x\,\partial_{r}\langle b^{-1}Ab^{\prime}\rangle-\int\operatorname{d}\!^{3}x\,\partial_{u}\langle\partial_{r}bb^{-1}g^{-1}g^{\prime}\rangle\right)+I_{\text{\tiny bdy}}\,.$ (2.19)
The last term in the parenthesis of(2.19) is zero if we take $u$ periodic,
@@
$I_{\text{\tiny CS}}$ $=\frac{k}{4\pi}\left(-\Gamma[G]+\int_{\partial\mathcal{M}}\langle g^{-1}\dot{g}\,g^{-1}g^{\prime}\rangle\right)+I_{\text{\tiny bdy}}\,.$ (2.21)
@@
$\int_{\partial\mathcal{M}}\langle g^{-1}\dot{g}\,g^{-1}g^{\prime}\rangle$ $=-\int\operatorname{d}\!u\operatorname{d}\!x\Big{[}e^{\Phi}\left(\dot{V}X^{\prime}+\dot{X}V^{\prime}+\dot{Y}W^{\prime}+\dot{W}Y^{\prime}+\zeta(\dot{X}Y^{\prime}+\dot{Y}X^{\prime})\right)$
@@
$\Gamma[G]=-\int\operatorname{d}\!^{3}x\,\partial_{r}\Big{[}e^{\Phi}\left(\dot{V}X^{\prime}-\dot{X}V^{\prime}+\dot{Y}W^{\prime}-\dot{W}Y^{\prime}+\zeta(\dot{Y}X^{\prime}-\dot{X}Y^{\prime})\right)\Big{]}\,.$ (2.24)
@@
$I_{{\text{\tiny CS}}}=-\frac{k}{2\pi}\int\operatorname{d}\!u\operatorname{d}\!x\Big{[}e^{\Phi}\left(\dot{X}(V^{\prime}+\zeta Y^{\prime})+\dot{W}Y^{\prime}\right)-\frac{1}{4}(\dot{\Phi}\zeta^{\prime}+\dot{\zeta}\Phi^{\prime})\Big{]}+I_{{\text{\tiny bdy}}}\,.$ (2.25)
@@
$I[\Phi,\zeta]=\frac{k}{8\pi}\int\operatorname{d}\!u\operatorname{d}\!x\left(\dot{\Phi}\zeta^{\prime}+\dot{\zeta}\Phi^{\prime}-\dot{\Phi}\right)\,,$ (2.29)
@@ 3 Reparametrization theory
$a(u)=a_{0}\,\dot{h}-\frac{\ddot{h}}{\dot{h}}\,,\qquad\eta(u)=\eta_{0}\,\dot{h}-\dot{\alpha}\,.$ (3.34)
@@
$I[h,\alpha]$ $=\frac{k}{8\pi}\int\operatorname{d}\!u\operatorname{d}\!x\Big{[}2a_{0}\eta_{0}\,\dot{h}h^{\prime}-a_{0}(\dot{h}\alpha^{\prime}+\dot{\alpha}h^{\prime})-\eta_{0}\left(h^{\prime}\frac{\ddot{h}}{\dot{h}}+\dot{h}^{\prime}\right)+\dot{\alpha}\frac{\dot{h}^{\prime}}{\dot{h}}+\alpha^{\prime}\frac{\ddot{h}}{\dot{h}}-a_{0}\dot{h}+\frac{\ddot{h}}{\dot{h}}\Big{]}\,.$ (3.37)
@@
It is easy to verify that these equations reproduce equations (2.2). The action (3.37) can be considered as the two-dimensional warped-Schwarzian theory when the U(1) Kac-Moody level is zero. The one-dimensional theory was introduced in Afshar:2019tvp as the group action for twisted warped coadjoint orbits of 1D diffeomorphisms on the circle. The vanishing level case is especially interesting because one can not remove the mixing term between $\alpha$ and $f$ via field redefinition. The same happens in the warped Virasoro algebra;
@@ 4 Boosted Rindler-AdS
$A^{\text{\tiny bRAdS}}=A^{\text{\tiny bRflat}}+\frac{1}{\ell}\left(-\operatorname{d}\!x\,L_{0}+\frac{1}{2}\eta\operatorname{d}\!x\,L_{-1}+\operatorname{d}\!u\,L_{1}\right)\,,$ (4.45)
where $A^{\text{\tiny bRflat}}$ is given in (2.8) for $\ell\to\infty$. In our parametrization of the algebra, the isl$(2,\mathbb{R})$ deforms to the so(2,2)22(2,2)( 2 , 2 ) algebra (2.7) for finite $\ell$. For the asymptotic AdS space, we use the same Gauss decomposition as the flat case (2.22), but here we take care of the added commutator, $[M_{n},M_{m}]=\frac{1}{\ell^{2}}(n-m)L_{m+n}$. As in the previous case, we assume $2\pi L$-periodicity in $u$. Again, assuming no nontrivial holonomies on the $x$-constant disc, we solve the constraint $F_{ru}=0$ by, $A_{u}=G^{-1}\dot{G}$ where $G$ is the arbitrary single valued group element of SO(2,2)22(2,2)( 2 , 2 ) which is decomposed according to (2.22). A precise calculation without the use of any specific representation for the group element leads to,
@@
$\Gamma[G]=$ $\frac{1}{3}\int\operatorname{d}\!^{3}x\,\epsilon^{\mu\nu\rho}\partial_{\mu}\Big{(}e^{\Phi}(\ell\partial_{\nu}X\partial_{\rho}Y+\frac{1}{\ell}\partial_{\nu}W\partial_{\rho}V)\text{sinh}\frac{\zeta}{\ell}+e^{\Phi}(\partial_{\nu}W\partial_{\rho}Y+\partial_{\nu}X\partial_{\rho}V)\text{cosh}\frac{\zeta}{\ell}\Big{)}\,.$
@@
$\int_{\partial{\mathcal{M}}}\langle g^{-1}\dot{g}\,g^{-1}g^{\prime}\rangle=$ $-\int\operatorname{d}\!u\operatorname{d}\!x\,e^{\Phi}\Big{[}(\dot{X}V^{\prime}+\dot{V}X^{\prime}+\dot{W}Y^{\prime}+\dot{Y}W^{\prime})\text{cosh}(\frac{\zeta}{\ell})$
@@
$I_{{\text{\tiny CS}}}=$ $-\frac{k}{2\pi}\int\operatorname{d}\!u\operatorname{d}\!x\,\Big{[}e^{\Phi}\Big{(}(\dot{X}V^{\prime}+\dot{W}Y^{\prime})\text{cosh}(\frac{\zeta}{\ell})+(\ell\dot{X}Y^{\prime}+\frac{\dot{W}V^{\prime}}{\ell})\text{sinh}(\frac{\zeta}{\ell})\Big{)}-\frac{1}{4}(\dot{\Phi}\zeta^{\prime}+\dot{\zeta}\Phi^{\prime})+\frac{1}{4}\,a\Big{]}\,,$ (4.49)
where in the last term we evaluated the boundary action (2.10) using the fact that for boundary condition (4.45) we have $A_{u}\,A_{x}\right>=-\frac{1}{2}a$. Again the action (4.49) agrees with (2.25) after taking the $\ell\to\infty$ limit.
@@
$I[\Phi,\zeta]=-\frac{k}{4\pi}\int\operatorname{d}\!u\operatorname{d}\!x\,\Big{[}2\frac{V^{\prime}}{\ell}+2Y^{\prime}-\frac{1}{2}(\dot{\Phi}\zeta^{\prime}+\dot{\zeta}\Phi^{\prime})+\frac{1}{2}\,a\Big{]}\,,$ (4.52)
again, we assume the fields are zero or have the same value at $x=\pm\infty$, so the first two terms can be ignored.
@@
$\delta a=\dot{\epsilon}\,a+\epsilon\,\dot{a}-\dot{\epsilon}-2\dot{\alpha}/\ell\,,\qquad\delta T=2\dot{\epsilon}\,a+\epsilon\,\dot{a}+\dot{\alpha}\,a+\ddot{\alpha}\,,$ (4.53)
@@
$T_{0}\,\dot{h}^{2}+\left(\frac{\ddot{h}}{\dot{h}}-a_{0}\,\dot{h}\right)(\dot{\zeta}+\dot{\alpha})-(\ddot{\zeta}+\ddot{\alpha})-\frac{1}{\ell}(\dot{\zeta}+\dot{\alpha})^{2}=0\,.$ (4.58)
@@
$I[h,\alpha]=\frac{k}{8\pi}\int\operatorname{d}\!u\operatorname{d}\!x\Big{[}$ $2T_{0}\,\dot{h}h^{\prime}-a_{0}\left(\dot{h}\alpha^{\prime}+h^{\prime}\dot{\alpha}\right)+\dot{\alpha}\frac{\dot{h}^{\prime}}{\dot{h}}+\alpha^{\prime}\frac{\ddot{h}}{\dot{h}}-a_{0}\dot{h}+\frac{\ddot{h}}{\dot{h}}-\frac{2}{\ell}\dot{\alpha}$
@@
which reduces to the action (3.37) in the flat limit $\ell\to\infty$. This action can be considered the two-dimensional warped-Schwarzian theory when the U(1) Kac-Moody level is present. The field equations corresponding to variation of $\alpha$ and $h$ are the following

### v4

# The near horizon dynamics in three-dimensional Einstein gravity

@@ 1 Introduction
The black hole near horizon physics has recently gained new attention in the framework of Jackiw-Teitelboim (JT) gravity Jackiw:1984je ; Teitelboim:1983ux . JT gravity as a two-dimensional dilaton gravity describes the dynamics close to the horizon of near-extremal black holes. Under appropriate boundary conditions it provides a toy model for holography in the Rindler AdS2 background, sometimes referred to as near AdS2/near CFT1 correspondence Kitaev:15ur ; Kitaev:2017awl ; Almheiri:2014cka ; Maldacena:2016hyu ; Maldacena:2016upp ; Engelsoy:2016xyb ; Cvetic:2016eiv ; Jensen:2016pah . Its holographic description is in terms of a one-dimensional Schwarzian action at the boundary for the Euclidean time reparametrization (which is the element of the asymptotic symmetry group). Similar holographic correspondences have been scrutinized in the 2D flat Rindler background (a near horizon description of non-extremal black holes) for flat JT-models Callan:1992rs ; Cangemi:1992bj ; Afshar:2019axx ; Godet:2020xpk ; Afshar:2020dth ; Godet:2021cdl ; Afshar:2021qvi ; Afshar:2022mkf . The Euclidean boundary dynamics in all such cases is in terms of an effective 1D Schwarzian-like action for elements of the asymptotic symmetry groups Afshar:2019tvp ; Afshar:2021qvi .
@@
The structure of the paper is as follows, in section 2 after introducing the 3D Rindler gravity theory we offer a WZW formulation for the theory using angular quantization of the Chern-Simons formulation. Then via Gauss decomposition of the gauge group, we obtain a Liouville-like boundary action for the flat Rindler boundary condition. In section 3 we promote this theory to a reparametrization theory in terms of elements of the warped Virasoro group. In section 4 we develop the same discipline for Rindler AdS3 and finally we discuss various aspects of the holonomy conditions in section 5.
@@ 2 Rindler boundary conditions
We are fully off-shell unless stated explicitly in the rest of the paper. We emphasize here that both functions $a$ and $\eta$ in our setup are state-dependent and thus are allowed to vary. This feature is qualitatively (and quantitatively) different from the alternative near-horizon boundary conditions Donnay:2015abr ; Afshar:2016kjj ; Afshar:2016uax ; Donnay:2016iyk and the ultimate near-horizon dynamics Grumiller:2019tyl where Rindler acceleration function is held fixed. As a consequence of imposing this set of boundary conditions, the asymptotic form of the metric (2.1) is preserved by diffeomorphisms generated by a vector field $\xi$ producing a Lie derivative such that $\mathcal{L}_{\xi}g_{xr}={\mathcal{L}}_{\xi}g_{xx}={\mathcal{L}}_{\xi}g_{rr}={\mathcal{L}}_{\xi}g_{ru}=0$;
$\xi^{u}=\epsilon(u)+\mathcal{O}(1/r)\,,\qquad\xi^{x}=\alpha(u)+\mathcal{O}(1/r)\,,\quad\xi^{r}=-r\dot{\epsilon}(u)+\mathcal{O}(1)\,,$ (2.4)
@@
Asymptotic region.
The Penrose diagram of the Rindler spacetime corresponding to the zero mode solution in (2.1) with $a=a_{0}$, $\eta=0$ and $\ell\to\infty$ is shown in Fig. 1. The Killing vector $\partial_{u}$ should be timelike, thus we have $g_{uu}=-2a_{0}r<0$ which says that we either have $(r>0,a_{0}>0)$ or $(r<0,a_{0}<0)$ corresponding to the regions I and III respectively. We thus have two asymptotic regions in general. However, there are different choices for placing the $u,r$ coordinates to generate a Rindler wedge. Here, instead of considering this pair of Rindler wedges, the setup we are considering is the Rindler wedge region I as the portion of the Penrose diagram of flat spacetime depicted in Fig. 2 covered by $r>0$ (and thus $a_{0}>0$). Thus we are dealing with only one asymptotic region $r\to\infty$. The ($u,r)$ coordinates in the region I can be viewed as describing the exterior of a finite temperature black hole accessible by the near horizon observer Maldacena:2016upp .
Figure 1: Penrose diagram of the Rindler spacetime.
Figure 2: Rindler wedge portion of the Penrose diagram of flat spacetime.
@@ 2.1 Chern-Simons formulation
The action that gives 3D Einstein gravity $I=\frac{1}{16\pi G}\int\operatorname{d}\!^{3}x\sqrt{-g}(R\pm 2/\ell^{2})$ is the Chern-Simons action up to a possible boundary term, for the gauge group ISO$(2,1)$, SO(2,2)22(2,2)( 2 , 2 ) or SO$(3,1)$ depending on the value of the cosmological constant $\Lambda=\mp\frac{1}{\ell^{2}}$ Achucarro:1986vz ; Witten:1988hc ,
@@
where $L_{\pm 1}=J_{0}\pm J_{1}$, $M_{\pm 1}=P_{0}\pm P_{1}$ and $L_{0}=J_{2}$ and $M_{0}=P_{2}$ and thus $\langle L_{1}\,M_{-1}\rangle=\langle L_{-1}\,M_{1}\rangle=-2\langle L_{0}\,M_{0}\rangle=-2$. We can recast our boundary conditions (2.1) in the first order form in terms of an algebra-valued connection, [^1] We have made a finite gauge transformation with the SO(2,2)22(2,2)( 2 , 2 ) group element $g=e^{\frac{r}{2}(M_{-1}-\frac{1}{\ell}L_{-1})}$ to eliminate the $r$-dependence. All results will be valid in the flat limit $\ell\to\infty$ for the theory with gauge group ISO$(2,1)$.
@@
$\delta I_{\text{{\tiny CS}}}=-\frac{k}{4\pi}\int\operatorname{d}\!^{3}x\,\epsilon^{r\nu\rho}\partial_{r}\langle A_{\nu}\delta A_{\rho}\rangle+\delta I_{\text{{\tiny bdy}}}\,.$ (2.9)
@@
$\langle A_{u}A_{x}\rangle=a+\mathcal{O}(1/\ell)\,.$ (2.11)
@@ 2.2 From Chern-Simons to Wess-Zumino-Witten
where $x^{\alpha}=(r,u)$ are coordinates on the $x$-constant slices, the Chern-Simons action (2.5) becomes,
$I_{\text{\tiny{CS}}}=\frac{k}{4\pi}\int\operatorname{d}\!^{3}x\,\epsilon^{\alpha\beta}\left\langle-A_{\alpha}\partial_{x}A_{\beta}+A_{x}F_{\alpha\beta}\right\rangle-\frac{k}{4\pi}\int\operatorname{d}\!^{3}x\partial_{r}\langle A_{u}A_{x}\rangle+\frac{k}{4\pi}\int\operatorname{d}\!^{3}x\partial_{u}\langle A_{r}A_{x}\rangle+I_{\text{\tiny bdy}}\,,$ (2.13)
where $I_{\text{\tiny bdy}}$ is given in (2.9) and is equivalent with the first total $r$-derivative term appearing in (2.13), so they add up to $2I_{\text{\tiny bdy}}$. The total $u$-derivative term is zero for our boundary conditions. Here by assuming that there are no nontrivial holonomies on the $x$-constant plane, a pure gauge configuration solves the constraint $F_{ru}=0$ by,
$A_{\alpha}=G^{-1}\partial_{\alpha}G\,,\qquad\alpha=r,u\,,$ (2.14)
@@
$I_{\text{\tiny CS}}$ $=\frac{k}{4\pi}\int\operatorname{d}\!^{3}x\,\langle A_{u}A_{r}^{\prime}-A_{r}A_{u}^{\prime}\rangle+2I_{\text{\tiny bdy}}$
$=\frac{k}{4\pi}\int\operatorname{d}\!^{3}x\,\langle-b^{-1}Ab^{\prime}b^{-1}\partial_{r}b+b^{-1}A\partial_{r}b^{\prime}+\partial_{r}bb^{-1}b^{\prime}b^{-1}A-\partial_{r}bb^{-1}A^{\prime}-b^{-1}\partial_{r}bb^{-1}Ab^{\prime}\rangle+2I_{\text{\tiny bdy}}\,.$ (2.17)
@@
$I_{\text{\tiny CS}}$ $=\frac{k}{4\pi}\left(-\Gamma[G]+\int\operatorname{d}\!^{3}x\,\partial_{r}\langle b^{-1}Ab^{\prime}\rangle-\int\operatorname{d}\!^{3}x\,\partial_{u}\langle\partial_{r}bb^{-1}g^{-1}g^{\prime}\rangle\right)+2I_{\text{\tiny bdy}}\,.$ (2.19)
The last term in the parenthesis of (2.19) is zero if we take $u$ to be periodic,
@@
$I_{\text{\tiny CS}}$ $=\frac{k}{4\pi}\left(-\Gamma[G]+\int_{\partial\mathcal{M}}\langle g^{-1}\dot{g}\,g^{-1}g^{\prime}\rangle\right)+2I_{\text{\tiny bdy}}\,.$ (2.21)
@@
$\int_{\partial\mathcal{M}}\langle g^{-1}\dot{g}\,g^{-1}g^{\prime}\rangle$ $=-2\int\operatorname{d}\!u\operatorname{d}\!x\Big{[}e^{\Phi}\left(\dot{V}X^{\prime}+\dot{X}V^{\prime}+\dot{Y}W^{\prime}+\dot{W}Y^{\prime}+\zeta(\dot{X}Y^{\prime}+\dot{Y}X^{\prime})\right)$
@@
$\Gamma[G]=-2\int\operatorname{d}\!^{3}x\,\partial_{r}\Big{[}e^{\Phi}\left(\dot{V}X^{\prime}-\dot{X}V^{\prime}+\dot{Y}W^{\prime}-\dot{W}Y^{\prime}+\zeta(\dot{Y}X^{\prime}-\dot{X}Y^{\prime})\right)\Big{]}\,.$ (2.24)
@@
$I_{{\text{\tiny CS}}}=-\frac{k}{\pi}\int\operatorname{d}\!u\operatorname{d}\!x\Big{[}e^{\Phi}\left(\dot{X}(V^{\prime}+\zeta Y^{\prime})+\dot{W}Y^{\prime}\right)-\frac{1}{4}(\dot{\Phi}\zeta^{\prime}+\dot{\zeta}\Phi^{\prime})\Big{]}+2I_{{\text{\tiny bdy}}}\,.$ (2.25)
@@
$I_{\text{\tiny RFlat}}[\Phi,\zeta]=\frac{k}{4\pi}\int\operatorname{d}\!u\operatorname{d}\!x\left(\dot{\Phi}\zeta^{\prime}+\dot{\zeta}\Phi^{\prime}-2\dot{\Phi}\right)\,,$ (2.29)
@@
We emphasize that (2.29) is the boundary action obtained upon reducing gravity in region I (see fig. 2) to its sole right asymptotic boundary as a consequence of restricting ourselves to the right Rindler wedge as explained in the beginning of section 2. This choice is valid for us because our goal is ultimately to use analytic continuation to translate our asymptotic dynamics in favor of the Euclidean theory in section 5 which possesses only one boundary. In general, if one considers both asymptotic regions in the Lorentzian theory, reducing the gravity theory to boundary dynamics is more complicated as one should include Wilson lines stretched between left and right boundaries. For consulting details of this line of analysis follow Harlow:2018tqv ; Cotler:2020ugk ; Henneaux:2019sjx .
@@ 3 Reparametrization theory
$a(u)=a_{0}\,\dot{h}-\frac{\ddot{h}}{\dot{h}}\,,\qquad\eta(u)=\eta_{0}\,\dot{h}-\dot{\alpha}\,,$ (3.34)
@@
$I[h,\alpha]$ $=\frac{k}{4\pi}\int\operatorname{d}\!u\operatorname{d}\!x\Big{[}2a_{0}\eta_{0}\,\dot{h}h^{\prime}-a_{0}(\dot{h}\alpha^{\prime}+\dot{\alpha}h^{\prime})-\eta_{0}\left(h^{\prime}\frac{\ddot{h}}{\dot{h}}+\dot{h}^{\prime}\right)+\dot{\alpha}\frac{\dot{h}^{\prime}}{\dot{h}}+\alpha^{\prime}\frac{\ddot{h}}{\dot{h}}-2a_{0}\dot{h}+2\frac{\ddot{h}}{\dot{h}}\Big{]}\,.$ (3.37)
@@
It is easy to verify that these equations reproduce equations (2.2). The action (3.37) can be considered as the two-dimensional warped-Schwarzian theory when the U(1) Kac-Moody level is zero. The one-dimensional theory was introduced in Afshar:2019tvp as the group action for the twisted warped Virasoro coadjoint orbits on the circle. The vanishing level case is especially interesting because one can not remove the mixing term between $\alpha$ and $h$ via field redefinition. The same happens in the warped Virasoro symmetry algebra;
@@ 4 Boosted Rindler-AdS
$A^{\text{\tiny RAdS}}=A^{\text{\tiny Rflat}}+\frac{1}{\ell}\left(-\operatorname{d}\!x\,L_{0}+\frac{1}{2}\eta\operatorname{d}\!x\,L_{-1}+\operatorname{d}\!u\,L_{1}\right)\,,$ (4.45)
where $A^{\text{\tiny Rflat}}$ is given in (2.8) for $\ell\to\infty$. In our parametrization of the algebra, the isl$(2,\mathbb{R})$ deforms to the so(2,2)22(2,2)( 2 , 2 ) algebra (2.7) for finite $\ell$. For the asymptotic AdS space, we use the same Gauss decomposition as the flat case (2.22), but here we take care of the added commutator, $[M_{n},M_{m}]=\frac{1}{\ell^{2}}(n-m)L_{m+n}$. As in the previous case, we assume $2\pi L$-periodicity in $u$. Again, assuming no nontrivial holonomies on the $x$-constant disc, we solve the constraint $F_{ru}=0$ by, $A_{u}=G^{-1}\dot{G}$ where $G$ is the arbitrary single valued group element of SO(2,2)22(2,2)( 2 , 2 ) which is decomposed according to (2.22). A precise calculation without the use of any specific representation for the group element leads to,
@@
$\Gamma[G]=$ $\frac{2}{3}\int\operatorname{d}\!^{3}x\,\epsilon^{\mu\nu\rho}\partial_{\mu}\Big{(}e^{\Phi}(\ell\partial_{\nu}X\partial_{\rho}Y+\frac{1}{\ell}\partial_{\nu}W\partial_{\rho}V)\text{sinh}\frac{\zeta}{\ell}+e^{\Phi}(\partial_{\nu}W\partial_{\rho}Y+\partial_{\nu}X\partial_{\rho}V)\text{cosh}\frac{\zeta}{\ell}\Big{)}\,.$
@@
$\int_{\partial{\mathcal{M}}}\langle g^{-1}\dot{g}\,g^{-1}g^{\prime}\rangle=$ $-2\int\operatorname{d}\!u\operatorname{d}\!x\,e^{\Phi}\Big{[}(\dot{X}V^{\prime}+\dot{V}X^{\prime}+\dot{W}Y^{\prime}+\dot{Y}W^{\prime})\text{cosh}(\frac{\zeta}{\ell})$
@@
$I_{{\text{\tiny CS}}}=$ $-\frac{k}{\pi}\int\operatorname{d}\!u\operatorname{d}\!x\,\Big{[}e^{\Phi}\Big{(}(\dot{X}V^{\prime}+\dot{W}Y^{\prime})\text{cosh}(\frac{\zeta}{\ell})+(\ell\dot{X}Y^{\prime}+\frac{\dot{W}V^{\prime}}{\ell})\text{sinh}(\frac{\zeta}{\ell})\Big{)}-\frac{1}{4}(\dot{\Phi}\zeta^{\prime}+\dot{\zeta}\Phi^{\prime})+\frac{1}{2}\,a\Big{]}\,,$ (4.49)
where in the last term we evaluated the boundary action (2.10) using the fact that for boundary condition (4.45) we have $A_{u}\,A_{x}\right>=a$. Again the action (4.49) agrees with (2.25) after taking the $\ell\to\infty$ limit.
@@
$I_{\text{\tiny RAdS}}[\Phi,\zeta]=-\frac{k}{2\pi}\int\operatorname{d}\!u\operatorname{d}\!x\,\Big{[}2\frac{V^{\prime}}{\ell}+2Y^{\prime}-\frac{1}{2}(\dot{\Phi}\zeta^{\prime}+\dot{\zeta}\Phi^{\prime})+a\Big{]}\,,$ (4.52)
again, we assume the fields are zero or have the same value at $x=\pm\infty$, so the first two terms can be ignored and thus we get $I_{\text{\tiny RAdS}}=I_{\text{\tiny RFlat}}+\frac{k}{2\pi\ell}\int\operatorname{d}\!^{2}x\,\dot{\zeta}$, where $I_{\text{\tiny RFlat}}$ is given in (2.29).
@@
$\delta a=\dot{\epsilon}\,a+\epsilon\,\dot{a}-\ddot{\epsilon}-2\dot{\alpha}/\ell\,,\qquad\delta T=2\dot{\epsilon}\,a+\epsilon\,\dot{a}+\dot{\alpha}\,a+\ddot{\alpha}\,,$ (4.53)
@@
$T_{0}\,\dot{h}^{2}+\left(\frac{\ddot{h}}{\dot{h}}-a_{0}\,\dot{h}\right)(\dot{\zeta}+\dot{\alpha})-(\ddot{\zeta}+\ddot{\alpha})-\frac{1}{\ell}(\dot{\zeta}+\dot{\alpha})^{2}=0\,,$ (4.58)
@@
$I[h,\alpha]=\frac{k}{4\pi}\int\operatorname{d}\!u\operatorname{d}\!x\Big{[}$ $2T_{0}\,\dot{h}h^{\prime}-a_{0}\left(\dot{h}\alpha^{\prime}+h^{\prime}\dot{\alpha}\right)+\dot{\alpha}\frac{\dot{h}^{\prime}}{\dot{h}}+\alpha^{\prime}\frac{\ddot{h}}{\dot{h}}-2a_{0}\dot{h}+2\frac{\ddot{h}}{\dot{h}}-\frac{4}{\ell}\dot{\alpha}$
@@
which reduces to the action (3.37) in the flat limit $\ell\to\infty$. This action can be considered as a two-dimensional warped-Schwarzian theory when the U(1) Kac-Moody level is present. The field equations corresponding to variation of $\alpha$ and $h$ are the following